# Classical simulability of constant-depth linear-optical circuits with noise

### v1

@@ I Introduction
Quantum optical platforms using photons are expected to play versatile roles in various quantum information processing tasks, such as quantum communication, quantum sensing, and quantum computing [1, 2, 3, 4, 5, 6]. Especially, quantum optical circuits using linear optics are more experimentally feasible and still have the potential to provide a quantum advantage for quantum computing; representative examples are boson sampling using single photons [7] or Knill-Laflamme-Milburn (KLM) protocol for universal quantum computation [8].
@@
However, as in other experimental platforms, one of the main obstacles to implementing a large-scale quantum device to perform interesting quantum information processing is noise; especially, photon loss and partial distinguishability of photons in photonic devices are typically the most crucial noise sources [9, 10, 11, 12]. Through experimental realizations of intermediate-scale quantum devices using photons and thorough theoretical analysis of the effect of loss and noise, many recent results show that they can significantly reduce the computational power of the quantum devices [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Hence, there have been significant interests and efforts in reducing the effect of photon loss and partial distinguishability, such as developing quantum error correction codes [22, 23, 24, 25, 26].
@@
Since the photon loss effect becomes more severe as circuit depth increases and eventually quantum circuits become easy to classically simulate in many cases [15, 16, 17, 19, 20, 21], for quantum computing, more specifically for demonstrating quantum computational advantage, another viable and promising path is to employ a shallow-depth quantum circuit to minimize the effect of photon loss. In fact, a worst-case constant-depth linear-optical circuit with single photons is proven to be hard to simulate exactly using classical computers unless the polynomial hierarchy (PH) collapses to a finite level [27]. Furthermore, there have been many attempts to prove the average-case hardness of approximate simulation of shallow-depth boson sampling circuits [28, 29, 30]. Since one of the main reasons to utilize shallow-depth circuits is to minimize the effect of noise and loss, a pertinent question that needs to be answered is whether shallow-depth circuits under the effect are hard to classically simulate or the effect again destroys the potential quantum advantage even from shallow-depth circuits.
@@
To address this question, in this work, we analyze the computational complexity of constant-depth linear-optical circuits under photon loss and partial distinguishability and prove that when the input state is single photons, there exists a threshold of noise rates above which we can efficiently simulate the system using classical computers. The main idea is to associate a linear-optical circuit with a bipartite graph in such a way that the single photons correspond to one part of the vertices of the graph and the output modes correspond to the other part of the vertices of the graph and they are connected by edges if the single photons can propagate to the output modes through the linear-optical circuit. We then appropriately adapt a well-known result of percolation theory from the study of network [31, 32] to bipartite graphs, which states that if some of the vertices of a graph of bounded degree are randomly removed, the resultant graph is divided into disjoint logarithmically-small-size graphs with high probability. We then show that the effect of photon loss or partial distinguishability noise exactly corresponds to removing some of the vertices; thus, photon loss or partial distinguishability of photons effectively transforms constant-depth linear-optical circuits into logarithmically-small-size independent linear-optical circuits. Consequently, we can simulate the entire circuit by individually simulating each small-size linear-optical circuit. Our result suggests that while shallow-depth circuits are often believed to be less subject to loss and noise, it may not always be true because the entanglement constituted by shallow-depth circuits may be more easily destroyed by loss and noise. Finally, we numerically analyze the effect for various architectures and discuss a general condition for our result to hold.
II Linear-optical circuits with single photons
@@
Let us consider $M$-mode linear-optical circuits with $N$ single photons as an input state and arbitrary local measurement for the output state. Here, linear-optical circuits are composed of layers of beam splitters, which may be geometrically non-local. This setup is the basis of single-photon boson sampling [7] or the KLM protocol [8] for universal quantum computation. While the latter typically require sufficiently deep linear-optical circuits, in this work, we will mainly focus on much shallower depth circuits, more precisely, constant-depth circuits, which are expected to be less subject to loss and noise. We emphasize again that constant-depth boson sampling is proven to be hard to classically simulate unless the PH collapses to a finite level [27]; thus, shallow-depth linear-optical circuits may be sufficient for quantum computational advantage.
@@
Figure 1: (a) Linear-optical circuit with input photons on red dots and vacuum on the empty dots on the left and output modes on blue dots on the right. Input and output modes are coupled by beam splitter arrays, which make the input photons propagate through the dotted lines. (b) The corresponding bipartite graph. (c) When the second and third photons get lost, the corresponding vertices are removed from the bipartite graph. Consequently, the resultant bipartite graph is divided into two disconnected bipartite graphs.
@@
III Bipartite-Graph percolation
Percolation theory describes the behavior of graphs obtained by adding or deleting vertices or edges of graphs [31, 32]. Percolation theory has also been used in quantum information theory to study entanglement in quantum networks [33, 34, 35]. Also, very recently, it has been used to invent a classical algorithm for simulating constant-depth IQP circuits [36]. Using a similar technique, we will show that constant-depth linear-optical circuits with single photons are easy to classically simulate when the loss rate or the noise rate is sufficiently high. The key idea for this, together with the percolation lemma below, is that if a single photon is lost or becomes distinguishable from others, the system then essentially loses interference induced from the photon. This effect, from graph theoretically point of view, is to remove the vertex or to decouple the vertex from the graph for loss and distinguishability noise, respectively (see below for more details). To use this property, we adapt a result of percolation theory from Refs. [37, 38, 36] to bipartite graphs (see Appendix A for the proof):
@@ III Bipartite-Graph percolation
Let $G=(A,B,E)$ be a bipartite graph of maximum degree $\Delta$. If we independently remove each $v\in A$ with probability $1-\eta$ with $\eta<1/\Delta^{2}$ and all edges incident to $v$ from $E$, then the resultant bipartite graph is divided into $m$ bipartite graphs $\{G_{i}\}_{i=1}^{m}$ disconnected to each other and
@@
From the physical point of view, the lemma implies that for a linear-optical circuit with maximum degree $\Delta$, losing ${1-1/\Delta^{2}}$ portion of input states is so significant that the remaining system can be effectively described by many independent small-size systems.
It is worth emphasizing that the above lemma can be understood from a previous result [38, 36] by defining a graph based on the bipartite graph in such a way that the graph’s vertices are $A$ and two vertices have an edge if they are connected by a vertex in $B$ in the original bipartite graph. Physically, two input modes are connected if the photons from them can be outputted in the same output mode, i.e., they can interfere each other. Using the relation between the underlying bipartite and the new graph $G$, we can easily see that the maximum degree of the graph $G$ is upper-bounded by $\Delta^{2}$.
IV Photon loss effect
Now, let us consider photon loss on input photons and observe the correspondence between photon loss and the removal of some vertices in the associated bipartite graph as in Lemma 1. When we prepare a single photon and then it is subject to a loss channel with loss rate $1-\eta$, the single-photon state transforms as
@@ IV Photon loss effect
Then, while there is no effect for the photon with probability $\eta$, i.e., the vertex of the associated bipartite graph is kept, the vertex is removed from $A$ with probability $1-\eta$. Consequently, when we prepare $N$ single photons and the single photons are subject to a loss channel with loss rate $1-\eta$, the effect is to remove each vertex with probability $1-\eta$ independently, which is exactly the same procedure in Lemma 1 (see Fig. 1 for the illustration).
By taking advantage of this relation and Lemma 1, we propose a classical algorithm simulating linear-optical systems with $N$ single photon input when $\eta\Delta^{2}<1$. First, we remove each vertex in $A$ with probability $1-\eta$ for the associated bipartite graph with a given linear-optical circuit with single photons. For the resultant graph, we identify components $\{G_{i}\}_{i=1}^{m}$ that are disconnected from each other. If the size of any connected component $G_{i}$ obtained from the first step is larger than $y^{*}$, then we return to the first step. Otherwise, we now have $\{G_{i}\}_{i=1}^{m}$ with $|G_{i}|\leq y^{*}=O(\log(N/\epsilon))$ for all $i$’s. Since the number of photons for each connected component scales logarithmically, we can expect that it is easy to classically simulate. Here, $\epsilon$ is chosen to be the desired total variation distance (see below).
More specifically, to see how we simulate each system associated with $G_{i}$ as desired since the number of input photons in each $G_{i}$ is at most $y^{*}$ and the number of relevant output modes is at most $\Delta y^{*}$, the associated Hilbert space’s dimension is upper-bounded by
@@
where we used $\binom{n}{k}\leq(\frac{ne}{k})^{k}$. Thus, when $\Delta=O(1)$, which is the case for a constant depth circuit, the dimension is upper bounded by $\text{poly}(N/\epsilon)$. Therefore, since writing down the output states and the relevant operators takes polynomial time in $N$ and $1/\epsilon$ for any local measurement, we can efficiently simulate the system. If the measurement is photon number detection, which corresponds to boson sampling [7], then one may simply use the Clifford-Clifford algorithm whose complexity is given by $\tilde{O}(2^{y^{*}})=\text{poly}(N/\epsilon)$ [39].
One may notice that if $\Delta$ scales with the system size $N$ and the measurement is not on a photon number basis, the above counting gives us a superpolynomially increasing dimension in $N/\epsilon$. For this case, to be more efficient, consider a $y^{*}$ number of single photons as an input and note that a linear-optical circuit $\hat{U}$ transforms the creation operators of input modes $\hat{a}^{\dagger}_{j}$ as
@@
Therefore, for any bipartition, the output state can be described by at most $2^{y^{*}}=\text{poly}(N/\epsilon)$ singular values, which means that there exists a matrix product state that can describe the state with bond dimension $\text{poly}(N/\epsilon)$ [40], which can be found by using time-evolution block decimation [41]. It is worth emphasizing that if the circuit is not linear-optical, the output bosonic operators may not be decomposed by a similar way and thus require more computational costs because it may contain an operator that is a product of two operators from each partition (e.g., $\hat{a}_{1}^{\dagger}\hat{a}_{M}^{\dagger}$).
The remaining question is the algorithm’s error, which is caused by skipping the case where the connected component’s size is larger than $y^{*}$. Recall that if we encounter the case where the maximum size of the connected components is larger than $y^{*}$, then we restart the sampling. Then, the output probability $q$ of such an algorithm is given by
@@
where $p(\bm{m}|E)$ is the conditional probability on the case $E$ where the maximum size of the connected components is smaller than or equal to $y^{*}$; hence, the true probability of the linear-optical circuit is written as
@@
where $E^{\perp}$ is the case where the maximum size of the connected components is larger than $y^{*}$. Then, the total variation distance between the suggested algorithm’s output probability $q(\bm{m})$ and the true output probability $p(\bm{m})$ can be shown to be smaller than $\epsilon$ (see Appendix C):
$\text{TVD}=\frac{1}{2}\sum_{\bm{m}}|p(\bm{m})-q(\bm{m})|\leq\epsilon.$ (10)
Finally, there is an overhead per sample due to the restart, but this is only $1/p(E)\leq 1/(1-\epsilon)=O(1)$, which is negligible. Thus, we obtain the following theorem:
@@
For a given loss rate $1-\eta$ and a linear-optical circuit of an arbitrary architecture of maximum degree $\Delta$ with $N$ single photon input, if $\eta<1/\Delta^{2}$, there exists a classical algorithm that can approximately simulate the corresponding lossy linear-optical circuits in $\text{poly}(N,1/\epsilon)$ within total variation distance $\epsilon$.
Hence, for constant-depth linear-optical circuits, i.e., $d=O(1)$ and thus $\Delta=O(1)$, there is a threshold of loss rate above which it becomes classically easy to sample. Note that although we assumed the maximum degree $\Delta$ to be $2^{d}$ to cover the worst case, the actual loss threshold depends on the architecture (see below for further discussion on this). It is worth emphasizing that the notion of approximate simulation is crucial because the exact classical simulation of constant-depth lossy linear-optical circuits is hard unless the PH collapses to a finite level. This can easily be shown by noting that postselecting no loss case of lossy boson sampling is equivalent to lossless boson sampling and constant-depth boson sampling with post-selection is post-BQP due to the measurement-based quantum computing [27]. Thus, if lossy constant-depth boson sampling can be exactly simulated using classical algorithms efficiently, $\text{PH}\subset\text{P}^{\text{PP}}=\text{P}^{\text{post-BQP}}=\text{P}^{\text{post-BPP}}$ [42], which contradicts the fact that $\text{P}^{\text{post-BPP}}$ is in the PH [43] assuming that the PH is infinite.
Here, we also consider cases where each layer of beam splitters has a transmission rate $\eta_{1}$ that is smaller than $1$, and thus $\eta=\eta_{1}^{d}$. Since $\Delta\leq 2^{d}$ for any architecture and loss channel with a loss rate commutes with beam splitters (see e.g., Refs. [15, 17]), we have the following corollary:
@@
For a given transmission rate per layer $\eta_{1}$ and a linear-optical circuit of an arbitrary architecture of maximum degree $\Delta$ with $N$ single photon input, if $\eta_{1}<1/4$, there exists a classical algorithm that can approximately simulate the corresponding lossy linear-optical circuits in $\text{poly}(N,1/\epsilon)$ within total variation distance $\epsilon$.
For this case, the presented matrix product state method is crucial because the depth may not be constant.
V Partial distinguishability noise

### v2

@@ I Introduction
Quantum optical platforms using photons are expected to play versatile roles in quantum information processing, such as quantum communication, sensing, and computing [1, 2, 3, 4, 5, 6]. Especially, quantum optical circuits using linear optics are more experimentally feasible and still have the potential to provide a quantum advantage for quantum computing; representative examples are boson sampling [7] or Knill-Laflamme-Milburn (KLM) protocol for universal quantum computation [8].
@@
However, as in other experimental platforms, one of the main obstacles to implementing a large-scale quantum device to perform interesting quantum information processing is noise; especially, photon loss and partial distinguishability of photons in photonic devices are typically the most crucial noise sources [9, 10, 11, 12]. Through experimental realizations of intermediate-scale quantum devices using photons and thorough theoretical analysis of the effect of loss and noise, many recent results show that they can significantly reduce the computational power of the quantum devices [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Hence, there have been significant interests and efforts in reducing the effect of photon loss and partial distinguishability, such as developing quantum error correction codes [24, 25, 26, 27, 28].
The key idea of most of the existing classical algorithms for simulating lossy systems is that their output state becomes classical when the loss rate increases as the system size grows [13, 16, 17, 29, 19, 21, 22, 23, 30]. In particular, Refs. [13, 16, 17, 29, 23] show that when the total transmission rate scales as $O(1/\sqrt{N})$, where $N$ is the input photon number, the circuit can be approximately described by a nonnegative quasi-probability distribution [31, 32] and thus be classically simulated efficiently. Also, other classical algorithms considering various types of noises and using an approximation of the output probability by low-degree polynomials from Refs. [33, 14, 15, 21] assume Haar-random linear-optical circuits for their algorithms to be guaranteed to be efficient, which inevitably requires deep circuits [34] (See Supplemental Material (SM) [35] for more detailed discussion.).
@@
Hence, another promising path to quantum advantage is to employ shallow-depth quantum circuits to minimize the photon-loss effect, where the loss rate does not necessarily increase with system size and the existing methods do not directly apply. In fact, a worst-case constant-depth linear-optical circuit with single photons is proven hard for classical computers to exactly simulate unless the polynomial hierarchy (PH) collapses to a finite level [36]. Furthermore, there have been many attempts to prove the average-case hardness of approximate simulation of shallow-depth boson sampling circuits [37, 38, 39]. In addition, there have been proposals and proof-of-principle experiments that implement linear-optical circuits with high-degree connectivity, or even all-to-all connectivity [40, 41, 42, 43, 44, 45]. Since one main reason to utilize shallow-depth circuits is to minimize the effect of noise and loss, a pertinent question that must be answered is whether shallow-depth circuits, even under the effect, are hard to classically simulate or it again destroys the potential quantum advantage. It is worth emphasizing that since lossy boson sampling is hard to exactly classically simulate unless the PH collapses to a finite level (see the main text for more details.), an appropriate notion of simulation here is not an exact but approximate simulation, which is also a usual notion of quantum advantage [7].
@@
To address this question, in this work, we analyze the computational complexity of constant-depth linear-optical circuits with all-to-all connectivity under photon loss and partial distinguishability and prove that when the input state is single photons, there exists a threshold of noise rates above which we can approximately simulate the system efficiently using classical computers. The main idea is to associate a linear-optical circuit with a bipartite graph in such a way that the single photons correspond to one part of the vertices of the graph and the output modes correspond to the other part of the vertices of the graph and they are connected by edges if the single photons can propagate to the output modes through the linear-optical circuit. We then appropriately adapt a well-known result of percolation theory from the study of network [46, 47] to bipartite graphs, which states that if some of the vertices of a graph of bounded degree are randomly removed, the resultant graph is divided into disjoint logarithmically-small-size graphs with high probability. We then show that the effect of photon loss or partial distinguishability noise exactly corresponds to removing some of the vertices; thus, photon loss or partial distinguishability of photons effectively transforms constant-depth linear-optical circuits into logarithmically-small-size independent linear-optical circuits. Consequently, we can simulate the entire circuit by individually simulating each small-size linear-optical circuit. Our result suggests that while shallow-depth circuits are often believed to be less subject to loss and noise, it may not always be true because the entanglement constituted by shallow-depth circuits may be more easily destroyed by loss and noise. Finally, we numerically analyze the effect for various architectures and discuss a general condition for our result to hold.
II Results
II.1 Linear-optical circuits with single photons
@@
Consider $M$-mode linear-optical circuits with $N$ single-photon input and arbitrary local measurement. Linear-optical circuits consist of beam-splitter layers, which may be geometrically non-local, i.e., all-to-all connectivity. This setup is the basis of boson sampling [7] or the KLM protocol [8]. While the latter typically require deep linear-optical circuits, we mainly focus on constant-depth circuits which are expected to be less subject to loss and noise. Here, we define depth as the number of layers of beam splitters that can be implemented in parallel. We emphasize that constant-depth boson sampling is proven hard to exactly classically simulate unless the PH collapses to a finite level [36]; thus, shallow-depth linear-optical circuits may be sufficient for quantum advantage. Note, however, that it is not yet resolved whether constant-depth linear-optical circuits can yield quantum advantage for approximate simulation and that, as mentioned in the introduction, approximate simulation is the usual notation for claiming quantum advantage (e.g., Ref. [7]).
@@
Figure 1: (a) Linear-optical circuit with input photons on red dots and vacuum on empty dots on the left and output modes on blue dots on the right. Input and output modes are coupled by beam-splitter arrays with depth $d=3$, which make the input photons propagate through the dotted lines. (b) The corresponding bipartite graph. (c) When the second and third photons get lost, the corresponding vertices are removed from the graph. Consequently, the resultant bipartite graph is divided into two disconnected bipartite graphs.
@@
II.2 Bipartite-graph percolation
Percolation theory describes the behavior of graphs when adding or deleting vertices or edges of graphs [46, 47]. It has been used in quantum information theory to study entanglement in quantum networks [48, 49, 50] and classical simulation of noisy quantum circuits [51, 52, 53, 54], including a very recent result that presents an efficient classical algorithm simulating constant-depth noisy IQP circuits [54]. Using a similar technique, we show that constant-depth linear-optical circuits with single photons are easy to classically simulate when the loss or noise rate is sufficiently high. The key idea for this, together with the percolation lemma below, is that if a single photon is lost or becomes distinguishable from others, the system essentially loses interference induced by the photon. This effect, from a graph perspective, is to remove or decouple the vertex from the graph for loss and distinguishability noise, respectively (see below for more details). To use this property, we adapt a result of percolation theory from Refs. [55, 56, 54] to bipartite graphs (see Methods):
@@ II.2 Bipartite-graph percolation
Let $G=(A,B,E)$ be a bipartite graph of maximum degree $\Delta$. If we independently remove each $v\in A$ with probability $1-\eta$ with $\eta<1/\Delta^{2}$ and all edges incident to $v$ from $E$, the resultant bipartite graph is divided into $m$ bipartite graphs $\{G_{i}\}_{i=1}^{m}$ disconnected to each other and
@@
From the physical perspective, it implies that for linear-optical circuits with maximum degree $\Delta$, losing ${1-1/\Delta^{2}}$ portion of input photons is so significant that the remaining system can be effectively described by independent small-size systems.
II.3 Photon loss effect
Now, let us consider photon loss on input photons and observe the correspondence between photon loss and the removal of some vertices in the associated bipartite graph as in Lemma 1. When a single photon is subject to a loss channel with loss rate $1-\eta$, it transforms as
@@ II.3 Photon loss effect
Then, while there is no effect with probability $\eta$, the vertex is removed from $A$ with probability $1-\eta$. Consequently, when $N$ input single photons are subject to a loss channel with loss rate $1-\eta$, the effect is to remove each vertex with probability $1-\eta$ independently, which is exactly the same procedure in Lemma 1 (see Fig. 1 for the illustration).
By taking advantage of this relation and Lemma 1, we propose a classical algorithm simulating linear-optical systems with $N$ single photon input when $\eta\Delta^{2}<1$. First, we remove each vertex in $A$ with probability $1-\eta$ for the associated bipartite graph with a given linear-optical circuit with single photons. For the resultant graph, we identify components $\{G_{i}\}_{i=1}^{m}$ that are disconnected from each other. If the size of any connected component $G_{i}$ obtained from the first step is larger than $y^{*}$, then we return to the first step. Otherwise, we now have $\{G_{i}\}_{i=1}^{m}$ with $|G_{i}|\leq y^{*}=O(\log(N/\epsilon))$ for all $i$’s. Since the number of photons for each connected component scales logarithmically, we can expect that it is easy to classically simulate. Here, $\epsilon$ is chosen to be the desired total variation distance (TVD) (see below).
More specifically, since the input photon number in each $G_{i}$ is at most $y^{*}$ and the number of relevant output modes is at most $\Delta y^{*}$, the associated Hilbert space’s dimension is upper-bounded by
@@
where we used $\binom{n}{k}\leq(\frac{ne}{k})^{k}$. Thus, when $\Delta=O(1)$, like constant-depth circuits, the dimension is upper-bounded by $\text{poly}(N/\epsilon)$. Therefore, since writing down the output state and the relevant operators takes polynomial time in $N/\epsilon$ for any local measurement, we can efficiently simulate the system. For photon-number detection, like boson sampling [7], one may simply use the Clifford-Clifford algorithm whose complexity is given by $\tilde{O}(2^{y^{*}})=\text{poly}(N/\epsilon)$ [57].
If $\Delta$ scales with the system size $N$, i.e. beyond constant depth, and the measurement is not photon-number detection, the above counting gives us a superpolynomially increasing dimension in $N/\epsilon$. For this case, to be more efficient, consider a $y^{*}$ number of single photons as an input and note that a linear-optical circuit $\hat{U}$ transforms the creation operators of input modes $\hat{a}^{\dagger}_{j}$ as
@@
Therefore, for any bipartition, the output state can be described by at most $2^{y^{*}}=\text{poly}(N/\epsilon)$ singular values, implying that a matrix product state (MPS) with bond dimension $\text{poly}(N/\epsilon)$ can describe the state [58], which can be found by time-evolution block decimation [59]. Note that if the circuit is not linear-optical, the above decomposition is no longer valid because output bosonic operators may contain a product of two operators from each partition (e.g., $\hat{a}_{1}^{\dagger}\hat{a}_{M}^{\dagger}$).
The remaining question is the algorithm’s error caused by skipping the case where $\max_{i}|G_{i}|>y^{*}$. Then, the output probability $q$ of such an algorithm is given by
@@
where $p(\bm{m}|E)$ is the conditional probability on the case $E$ where $\max_{i}|G_{i}|\leq y^{*}$; hence, the true output probability is written as
@@
where $E^{\perp}$ is the case where $\max_{i}|G_{i}|>y^{*}$. Then, the TVD between $q(\bm{m})$ and $p(\bm{m})$ can be shown smaller than $\epsilon$ (see Methods). Finally, an overhead per sample from the restart is given by $1/p(E)\leq 1/(1-\epsilon)=O(1)$. Thus, we have
@@
For a given loss rate $1-\eta$ and a linear-optical circuit of maximum degree $\Delta$ with $N$ single-photon input, if $\eta<1/\Delta^{2}$, there exists a classical algorithm that can approximately simulate the circuit in $\text{poly}(N,1/\epsilon)$ within TVD $\epsilon$.
Hence, for constant-depth linear-optical circuits, $d=O(1)$ and $\Delta=O(1)$, there is a threshold of loss rate above which it becomes classically easy to simulate. Note that the actual threshold depends on the architecture (see below for further discussion). We emphasize that the measurement basis can be arbitrary as long as it is local; thus, our theorem can be employed beyond boson sampling.
Here, the notion of approximate simulation is crucial because the exact classical simulation of constant-depth lossy linear-optical circuits is hard unless the PH collapses to a finite level. This can easily be shown by noting that postselecting no loss case of lossy boson sampling is equivalent to lossless boson sampling and constant-depth boson sampling with post-selection is post-BQP due to the measurement-based quantum computing [36]. Thus, if lossy constant-depth boson sampling can be exactly simulated using classical algorithms efficiently, $\text{PH}\subset\text{P}^{\text{PP}}=\text{P}^{\text{post-BQP}}=\text{P}^{\text{post-BPP}}$ [60], which contradicts the fact that $\text{P}^{\text{post-BPP}}$ is in the PH [61] assuming that the PH is infinite. Hence, the complexity of exact simulation and approximate simulation of constant-depth lossy boson sampling has a gap when the loss rate is above the threshold.
We also consider cases where each layer of beam splitters has transmission rate $\eta_{1}<1$, thus $\eta=\eta_{1}^{d}$. Since $\Delta\leq 2^{d}$ for any architecture and loss channel with a uniform loss rate commutes with beam splitters (see, e.g., Refs. [16, 19]), we have the following corollary:
@@
For a linear-optical circuit with $N$ single-photon input and a transmission rate per layer $\eta_{1}$, if $\eta_{1}<1/4$, there exists a classical algorithm that can approximately simulate the circuit in $\text{poly}(N,1/\epsilon)$ within TVD $\epsilon$.
For this case, the presented MPS method is crucial because the depth may not be constant.
II.4 Partial distinguishability noise